OpenStudy (anonymous):

a cone has a volume of 98pi. every dimension is multiplied by a scale factor and the new cone has a volume of 6272pi. what was the scale factor

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{initial~~volume=98\pi\\~\\ \text{every dimension is multiplied by a scale factor}\\~\\ \text{cone has 3 dimensions therefore}\\~\\ \text{let the scale factor be }x^3\\~\\ 98x^3=6272\\~\\ x^3=\dfrac{6272}{98}\\~\\ x^3=64\\~\\ x=4 }\end{align}\)